OpenStudy (miss.rose):

Can someone help? I will medal and fan! Seven less than one-third of a number is 5. Put this in number form and solve it. Please Help!

OpenStudy (anonymous):

So you can make the equation: \[\frac{1}{3}x-7=5\]

OpenStudy (miss.rose):

I thought it was supposed to be x/3 - 7 = 5 Can you show me how to solve it?

OpenStudy (anonymous):

That is the same thing, as what I wrote. :) You can put the x up there, without it making a difference.

OpenStudy (anonymous):

So you have:\[\frac{x}{3}-7=5\] First, add 7 on each side: \[\frac{x}{3}=12\] Now multiply with 3 \[x=12*3=?\]

OpenStudy (miss.rose):

12 * 3 = 36 I will be right back. I have to go do something, so go hold that thought.

OpenStudy (anonymous):

Well there is nothing more to the question :) That is the answer, the number is 36.

OpenStudy (miss.rose):

Okay. Thank you!